# Lightweight Learning from Actuation-Space Demonstrations via Flow Matching for Whole-Body Soft Robotic Grasping

### v1

@@ III-B2 Model Architecture
IV Learning in Simulation
IV-A Simulation and Data Collection
In simulation, the manipulated object moves on a planar surface without rotation, and its state $o$ is represented by two translational degrees of freedom ($C=2$). The SpiRob is aligned parallel to the tabletop, ensuring its motion plane coincides with the surface. The object is modeled as a cylinder of 30 mm diameter, placed with its base on the table, as shown in Fig. 3. Physical parameters such as mass, stiffness, and damping follow the settings in [8].
@@ IV-A1 Training Region within the Workspace
where $(x_{0},z_{0})=(0,0)$, $r_{\mathrm{in}}=0.15\,\mathrm{m}$, $r_{\mathrm{out}}=0.35\,\mathrm{m}$, $x_{\min}=-0.25\,\mathrm{m}$, $x_{\max}=0.25\,\mathrm{m}$, $z_{\min}=-0.1\,\mathrm{m}$, and the exclusion zone $x_{\mathrm{exc}}^{\min}=-0.05\,\mathrm{m}$, $x_{\mathrm{exc}}^{\max}=0.05\,\mathrm{m}$. Within this workspace, a narrow training region (highlighted in green in Fig. 4) was selected for data collection, covering less than 8% of the total workspace area.
IV-A2 Data Collection
Control is realized in terms of tendon forces that drive the SpiRob’s motion. For each expert simulated demo, we record the continuous control function $U_{1}^{*}(t)$, $t\in[0,T]$, representing the tendon forces over time, together with the corresponding object state $o\in\mathbb{R}^{2}$ denoting the sampled placement. The discrete control sequence used for training is then obtained by applying the sampling operator $\mathcal{I}^{-1}[\cdot]$ to $U_{1}$, yielding $U_{1}=\mathcal{I}^{-1}[U_{1}^{*}(t)]\in\mathbb{R}^{N\times 2}$. Fig. 3 shows a representative expert grasping demo in the simulation, where the robot gradually approaches and encloses the object while $(U_{1},o)$ are recorded for subsequent data processing.
@@ IV-B Training and Inference Details
We collected 30 expert demonstrations in MuJoCo, each providing a continuous expert control function $U_{1}^{*}(t)=[F_{1}^{*},F_{2}^{*}$] with variable duration $T$ determined by the task completion time,where $F_{1}^{*}$ and $F_{2}^{*}$ denote the left and right tendon control forces. This lightweight learning setup uses a small amount of training data, covering less than 8% of the robot’s reachable workspace (Fig. 4). To obtain fixed-length training data, the sampling operator $\mathcal{I}^{-1}[\cdot]$ was applied to each $U_{1}^{*}(t)$, yielding a discrete control sequence $U_{1}=[F_{1},F_{2}]=\mathcal{I}^{-1}[U_{1}^{*}(t)]\in\mathbb{R}^{N\times 2}$, $N=80$ represents uniformly sampled timesteps across $[0,T]$. The model was trained for 50,000 iterations using the Adam optimizer [39] with a learning rate of $10^{-4}$ for Eq. 2, on one NVIDIA GeForce RTX 4090 GPU, requiring approximately six hours to complete.
During inference, the rectified flow model was sampled with 30 iteration steps $\tau$ (Fig. 2C), producing an inferred discrete expert-like control sequence $\hat{U}_{1}$, which is then interpolated using $\mathcal{I}[\cdot]$ to reconstruct the corresponding continuous expert-like control sequence $\hat{U}_{1}^{*}(t)=[\hat{F}_{1}^{*},\hat{F}_{2}^{*}]=\mathcal{I}[\hat{U}_{1}]$ for execution on the physical robotic platform.
@@ V Sim-to-real Experiments
We validate the proposed framework on a SpiRob by transferring the learned policies from simulation to the physical platform. This section first introduces the experimental setup and control-signal processing for sim-to-real transfer, followed by validations in three scenarios: (i) workspace generalization from sparse demonstrations, (ii) geometric adaptability to object size variations, and (iii) robust dynamic scalability.
@@ V-A Experimental Setup
Figure 7: Experimental results of workspace generalization from sparse demonstrations and geometric adaptability to object size variations
@@ V-A2 Force-to-Displacement Signal Processing for Sim-to-Real Transfer
On the physical platform, tendon displacement control is adopted because, generally speaking in robotics, displacement control provides a more user-friendly interface than direct force control [40]. For rigid robots, displacement control alone is inadequate for highly dynamic and uncertain interaction tasks, as it cannot explicitly regulate contact forces. In contrast, soft robots inherently exhibit embodied mechanical intelligence, allowing direct control in the actuation space, which enables adaptive and optimized interactions even under displacement control. This approach offers both ease of implementation and the ability to preserve the intrinsic adaptability of soft bodies. Therefore, the hardware executes control through tendon displacement, with motor encoders providing accurate position feedback. To ensure consistency, the inferred expert-like force trajectories $\hat{U}_{1}^{*}(t)$ are first applied in simulation to compute the corresponding tendon displacement profiles $\Delta\hat{l}(t)$ (Eq. (7), shown in Fig. 6B), which are then used as executable motor commands.
@@ V-B1 Workspace Generalization from Sparse Demonstrations
Using the proposed training framework, only $30$ demonstrations were collected within a small sampling region, covering less than 8% of the robot’s reachable workspace. Despite this limited and localized training data, the learned Rectified Flow policy was evaluated on $1000$ uniformly generated target placements across the full workspace, with successful and failed validations shown in different colors in Fig. 7A, achieving a grasping success rate of 97.5%. These results demonstrate that the proposed method enables strong generalization from sparse, small-range training data to diverse unseen positions in the entire workspace, and the grasping behaviors observed in simulation are consistently reproduced on the physical SpiRob platform (Fig. 7B).
V-B2 Geometric Adaptability to Object Size Variations
The Rectified Flow policy was trained exclusively on cylindrical objects of $30$ mm diameter. During evaluation, the same policy was applied to cylinders of 20 mm (-33%) and 40 mm (+33%) without retraining, and the SpiRob successfully completed the grasping tasks in all cases. These results demonstrate that the learned policy, combined with the robot’s embodied mechanical intelligence, enables robust adaptation to object size variations of approximately $\pm$33%. Representative grasping results are shown in Fig. 7C.
@@ V-B3 Robust Dynamic Scalability
The adaptability of the learned policy was further examined under different execution speeds. Starting from a reference displacement trajectory with duration $T$, the temporal axis was rescaled to produce faster or slower motions while maintaining the same displacement pattern (Fig. 8). The SpiRob successfully completed grasping with compressed executions up to $T/5$ and extended executions up to $2T$. Within this range, the grasping performance remained stable, indicating that the learned control pattern is robust to temporal scaling and can be directly adjusted by modifying the execution period $T$.
@@ VI Insights: Distributional Control and Actuation-Space Demonstration
Besides, the learned policy exhibits a distributional control pattern. Through the Rectified Flow model, the inferred $\hat{U}_{1}$ is inherently discrete and represents the distributional control samples produced by the model, while $\hat{U}_{1}^{*}(t)=[\hat{F}_{1}^{*},\hat{F}_{2}^{*}]=\mathcal{I}[\hat{U}_{1}]$ denotes the corresponding continuous distributional force control functions obtained through least-squares curve fitting respectively. As shown in Fig. 9, 100 sampled force control functions vary individually yet form a consistent distribution, with all executions converging to successful grasps. Rather than encoding a single trajectory, the policy represents a family of feasible control solutions, each providing a different realization through which the robot achieves the same task. This distributional property enhances robustness to uncertainty.
@@
Figure 9: Distributional sampling of control sequences.
@@
This study presented a lightweight framework for learning from actuation-space demonstrations, which couples embodied mechanical intelligence with distributional control to achieve robust whole-body grasping in soft robots. By employing the Rectified Flow model, the policy efficiently learned coordinated actuation patterns that exploit the inherent compliance and redundancy of the soft body, enabling adaptive manipulation across diverse conditions. Experiments conducted on a SpiRob demonstrated data-efficient learning from only a small number of demonstrations, strong geometric adaptation to object variations, and scalable dynamic performance under uncertainty. These results confirm that actuation-space learning effectively transforms passive flexibility into functional control intelligence, substantially reducing the reliance on centralized controller. Future work will extend this framework to multi-contact and dynamic manipulation scenarios, and explore hierarchical policy structures for scalable and autonomous soft robotic intelligence.

### v2

@@ IV Learning in Simulation
IV-A Data Collection in Simulation
In simulation, the manipulated object is fixed on a planar surface with its state $o\in\mathbb{R}^{2}$ representing the object’s 2D Cartesian coordinates. The SpiRob is aligned parallel to the tabletop. The object is modeled as a cylinder of 30 mm diameter, and placed on the table, as shown in Fig. 3. Physical parameters such as mass, stiffness, and damping follow the settings in [5].
@@ IV-A1 Training Region within the Workspace
where $(x_{0},z_{0})=(0,0)$, $r_{\mathrm{in}}=0.15\,\mathrm{m}$, $r_{\mathrm{out}}=0.35\,\mathrm{m}$, $x_{\min}=-0.25\,\mathrm{m}$, $x_{\max}=0.25\,\mathrm{m}$, $z_{\min}=-0.1\,\mathrm{m}$, and the exclusion zone $x_{\mathrm{exc}}^{\min}=-0.05\,\mathrm{m}$, $x_{\mathrm{exc}}^{\max}=0.05\,\mathrm{m}$. Within this workspace, a narrow training region (highlighted in green in Fig. 4), covering less than 8% of the total area, was selected based on the empirically derived feasible grasping workspace for the target object size, with sampling designed to ensure coverage and generalization within this feasible region [5].
IV-A2 Data Collection Strategy
Control is realized in terms of tendon forces that drive the SpiRob’s motion. Expert demonstrations in simulation are generated via human teleoperation: an operator interactively controls the two tendons in real time (based on empirical experience and the grasping insights reported in [5]). Concretely, during each rollout the operator adjusts the left/right tendon force commands to sequentially (i) uncurl and approach the target, and (ii) wrap and tighten to achieve enclosure, while observing the robot motion and contact state in the simulator.
For each expert simulated demo, we record the continuous control function $U_{1}^{*}(t)$, $t\in[0,T]$, representing the tendon forces over time, together with the corresponding object state $o\in\mathbb{R}^{2}$ denoting the sampled placement. The discrete control sequence used for training is obtained by applying the sampling operator $\mathcal{I}^{-1}[\cdot]$ to $U_{1}^{*}(t)$, yielding $U_{1}=\mathcal{I}^{-1}[U_{1}^{*}(t)]\in\mathbb{R}^{N\times 2}$, where $N$ is the number of uniformly sampled control steps over $[0,T]$. Fig. 3 shows a representative teleoperated grasping demo in simulation, where the robot gradually approaches and encloses the object while $(U_{1},o)$ are recorded for subsequent data processing.
@@ IV-B Training and Inference Details
We collected 30 expert demonstrations in MuJoCo, each providing a continuous expert control function $U_{1}^{*}(t)=[F_{1}^{*},F_{2}^{*}$] with variable duration $T$ determined by the task completion time,where $F_{1}^{*}$ and $F_{2}^{*}$ denote the left and right tendon control forces. This lightweight learning setup uses a small amount of training data, covering less than $8\%$ of the robot’s reachable workspace (Fig. 4). To obtain fixed-length training data, the sampling operator $\mathcal{I}^{-1}[\cdot]$ was applied to each $U_{1}^{*}(t)$, yielding a discrete control sequence $U_{1}=[F_{1},F_{2}]=\mathcal{I}^{-1}[U_{1}^{*}(t)]\in\mathbb{R}^{N\times 2}$, $N=80$ represents uniformly sampled timesteps across $[0,T]$. The model was trained for 50,00050{,}000 iterations using the Adam optimizer [37] with a learning rate of $10^{-4}$ for Eq. 2, on one NVIDIA GeForce RTX 4090 GPU, requiring approximately $6$ hours to complete.
During inference, the rectified flow model was sampled with $30$ iteration steps $\tau$ (Fig. 2C), producing an inferred discrete expert-like control sequence $\hat{U}_{1}$, which is then interpolated using $\mathcal{I}[\cdot]$ to reconstruct the corresponding continuous expert-like control sequence $\hat{U}_{1}^{*}(t)=[\hat{F}_{1}^{*},\hat{F}_{2}^{*}]=\mathcal{I}[\hat{U}_{1}]$ for execution on the simulated robot.
IV-C Simulation Validation
Using the proposed training framework, only 30 demonstrations within a small region (less than 8% of the reachable workspace) were used for training. The learned Rectified Flow policy was then evaluated on 1000 uniformly distributed targets across the full workspace, achieving a 97.597.5% grasping success rate in simulation (Fig. 7A). The six-layer Transformer, as shown in Fig. 2, contains approximately 11 M parameters and requires $0.25$ s per sequence for inference. A lightweight single-layer variant reduces the parameter count to 0.7 M, trains in about one hour, and achieves a $94\%$ success rate.
IV-D Comparison with Baselines in Simulation
To evaluate the effectiveness of the proposed RF model, we compare it with regression-based baselines for grasping tasks, including MLP, CNN, and Transformer, under the same sparse training setting of 10 demonstrations and identical training epochs. As shown in Table I, across 100 test trials, our generative- based method consistently outperforms all regression-based approaches, demonstrating superior extrapolation, particularly in challenging working zones with limited expert data coverage.
TABLE I: Success rate comparison with baselines in simulation.
Method MLP CNN Transformer Ours
Success rate 56% 68% 65% 94%
@@ V Sim-to-real Experiments
This section first introduces the experimental setup and control-signal processing for sim-to-real transfer, followed by validations in three scenarios: (i) workspace generalization from sparse demonstrations, (ii) geometric adaptability to object size variations, and (iii) robust dynamic scalability.
@@ V-A Experimental Setup
Figure 7: Sim-to-real results of workspace generalization from sparse demonstrations and geometric adaptability to object size variations
@@ V-A2 Force-to-Displacement Signal Processing for Sim-to-Real Transfer
On the physical platform, tendon displacement control is adopted, as displacement control generally provides a more user-friendly and reliable interface than direct force control in robotic systems [38]. For rigid robots, displacement control alone is often insufficient for highly dynamic and uncertain interaction tasks, since it cannot explicitly regulate contact forces. In contrast, soft robots inherently exhibit embodied mechanical intelligence through compliance and redundancy, allowing adaptive and stable interactions to emerge even under displacement-based actuation. As a result, direct control in the actuation space via tendon displacement offers both ease of implementation and the ability to preserve the intrinsic adaptability of soft bodies.
Accordingly, the hardware executes control through tendon displacement, with motor encoders providing accurate position feedback. To ensure consistency between simulation and hardware execution, the inferred expert-like force trajectories $\hat{U}_{1}^{*}(t)$ are first applied in simulation offline to compute the corresponding tendon displacement profiles $\Delta\hat{l}(t)$, defined as
$\small\Delta\hat{l}_{i}(t)=l_{i}(q(t))-l_{i}(q_{0}),\quad i=1,2,$ (10)
where $l_{i}(q(t))$ denotes the tendon length at configuration $q(t)$ and $q_{0}=\mathbf{0}$ is the undeformed reference configuration. Note that trained at a nominal friction coefficient of 0.10, the RF model maintains successful execution under ±50% friction mismatch in both simulation and hardware, indicating sufficient inherent generalization without domain randomization. The resulting displacement profiles $\Delta\hat{l}(t)$ are then directly used as executable motor commands on the physical platform (Fig. 6B).
@@ V-B Experimental Validation
Based on the simulation results, 50 representative targets were uniformly selected across the workspace for real-world experiments to evaluate the SpiRob’s workspace coverage, geometric adaptability, and dynamic scalability.
@@ V-B1 Workspace Generalization from Sparse Demonstrations
To verify real-world performance, the 50 representative targets were tested on the physical SpiRob platform, achieving a $100\%$ success rate and confirming strong generalization and consistent sim-to-real transfer (Fig. 7B).
V-B2 Geometric Adaptability to Shape and Size Variations
The Rectified Flow policy was trained exclusively on 30 mm cylinders. During evaluation, the same policy was applied to diverse geometries without retraining: cylinders ranging from 20 mm ($33\%$) to 60 mm ($100\%$), as well as triangular, square, hexagonal, and star-shaped prisms. The SpiRob successfully completed all grasping tasks across these variations. These results demonstrate that the learned policy, combined with the robot’s embodied mechanical intelligence, enables robust adaptation to both size variations ($33\%$ to $100\%$) and cross-shape generalization (Fig. 7C).
@@ V-B3 Robust Dynamic Scalability
The policy’s adaptability was further tested under different execution speeds. Starting from a reference trajectory with duration $T$, the temporal axis was rescaled to generate faster or slower motions while maintaining the same displacement pattern (Fig. 8). The SpiRob achieved stable grasping performance for speeds ranging from 5× faster ($T/5$) to 2× slower ($2T$) than the nominal duration, demonstrating robust temporal scalability of the learned control pattern.
@@ VI Insights: Distributional Control and Actuation-Space Demonstration
Besides, the learned policy via the RF model exhibits a distributional control pattern. As shown in Fig. 9, 100 repeated inferences at one fixed target position produce force control functions that vary individually yet form a consistent distribution, with all executions converging to successful grasps. Rather than encoding a single trajectory, the policy represents a family of feasible control solutions, each providing a different realization through which the robot achieves the same task. This distributional property enhances robustness to uncertainty.
@@
Figure 9: Distributional sampling of control sequences at one target position
@@
This study presents a lightweight framework for learning from actuation-space demonstrations that couples embodied mechanical intelligence with distributional control to enable robust whole-body grasping in soft robots. Using a Rectified Flow model, the policy learns coordinated actuation patterns that exploit the soft body’s compliance and redundancy. Experiments on a SpiRob show data-efficient learning from few demonstrations, strong generalization to object variations, and robust performance under uncertainty, reducing reliance on precise centralized control.The evaluation of this paper focuses on a simplified yet contact-rich setup: targets are constrained (fixed) on a horizontal plane and parameterized by their 2D positions. While this design enables repeatable benchmarking and isolates the grasp-synthesis problem from uncontrolled object motion, it does not capture the full complexity of free-object, 3D, multi-contact manipulation.Future work will extend this framework to free-object and dynamic multi-contact manipulation, incorporate richer task suites (e.g., grasp-and-place), explore hierarchical policy structures, and investigate co-design methods [39] that jointly optimize morphology (e.g., spiral geometry and tendon routing) and control policies.